Circular No. 7633

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bt IN IC 4830 R. Chassagne, Ste. Clotilde, Ile de Réunion, reports the discovery of

an apparent supernova (mag ∼ 15.0) on unfiltered CCD images taken with a 0.30-m f/2.7 telescope on May 24.0 and 24.9 UT. The object is located at α = 19^{h}13^{m}46^{s}.75, δ = −59^{o}17′22′′.8 (equinox 2000.0), which is 27′′.7 west and 17′′.1 north of the nucleus of IC 4830. An image of the same field taken on May 5.0 showed nothing at this position (limiting mag 19). The U.K. Schmidt image in the Digital Sky Survey (1975 May 10.7 UT) also showed nothing at this position (limiting mag 21).

1999 KW_{4}

P. Pravec and L. Šarounová, Ondřejov Observatory, report: “Our photometric observations of this object (cf. IAUC 7632) on six nights during 2000 June 19.0–27.9 UT revealed a generally nonperiodic lightcurve, the minima repeating on a characteristic interval of ∼ 1.5 hr with amplitudes 0.1–0.2 mag. This suggests that a slightly elongated primary rotates with a period of ∼ 3 hr, the deviations from strict periodicity perhaps being caused by the rotation lightcurve of and/or mutual occultations involving a secondary component (Pravec et al. 2000, Icarus 146, 190; see also Sinnott 2001, Sky Tel. 100, No. 6, 100). Extensive observations during this apparition are needed to describe the unusual photometric behavior and to estimate the parameters of the binary system.”

COMET C/1999 T2 (LINEAR) K. Hornoch, Lelekovice, Czech Republic, reports observations (0.35-m

reflector + CCD) of an antitail on several occasions during the past month: Apr. 24.89 UT, 4′.0 long in p.a. 51^{o}; May 10.92, 3′.0, 46^{o}; 20.95, 5′.0, 57^{o}; 24.94, 7′, 50^{o}. The main tail ranged from 1′.1 in p.a. 150^{o} on Apr. 24 to 1′.5 in p.a. 175^{o} on May 20; on May 24 there was a second broad tail 1′.8 long in p.a. 104^{o}, a very bright 0′.6 jet in p.a. 270^{o} and a much fainter 0′.5 jet in p.a. 200^{o}. CCD observations by J. Manteca, Barcelona, Spain (0.31-m Schmidt-Cassegrain), on May 23.00 put this comet at R = 15.0.

NOVA AQUILAE 2001 CCD photometry: May 20.07 UT, V = 13.33±0.08 (P. Sobotka, Brno,

Czech Republic; comparison GSC 1048.174: V = 14.52 by A. Henden); 23.07, 13.43±0.05 (Sobotka); 24.19, 13.42±0.1 (J. D. West, Mulvane, KS).

2001 May 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden